# Gravitational Effects of a Small Primordial Black Hole Passing Through the Human Body

### v1

@@ I Introduction
The detection of gravitational radiation from merging black holes LIGO as well as the direct imaging of M87* M87* and Sgr A* SgrA* has led to a significant resurgence of interest in black hole physics. Among the ideas that have been reexplored is the possibility that primordial black holes (PBH) could account for some or all of the dark matter in the Universe Carr(2020) . Many constraints can be placed on the mass of black holes that can serve as dark matter; the main open mass window lies in the asteroid mass range Carr(2021) ; Gorton(2024)
@@
Clearly, a sufficiently large PBH would have a significant impact on the human body; the goal of this present work is to determine the minimum black hole size needed to produce significant injury, and to see whether the observed lack of such injuries can place any bounds on PBH properties. Our discussion is partially motivated by earlier work on macroscopic dark matter (MACROs) Sidhu(2020) . In that paper, it was shown that the impact of MACRO dark matter on the human body would be sufficiently destructive that the non-observance of such impacts allows useful limits to be placed on the MACRO mass and cross section. While it is well known that a microscopic black hole would have a deleterious effect on the human body, there are no precise calculations. Most previous discussions of this topic are in the popular literature Loeb(2021) ; in fact, the first reference to the effect of a PBH on the human body appears to be in a work of fiction Niven(1974) .

### v2

@@ I Introduction
The detection of gravitational radiation from merging black holes LIGO as well as the direct imaging of M87* M87* and Sgr A* SgrA* has led to a significant resurgence of interest in black hole physics. Among the ideas that have been reexplored is the possibility that primordial black holes (PBH) could account for some or all of the dark matter in the Universe Carr(1974) ; Chapline(1975) ; Carr(1975) ; Khlopov(2010) ; Belotsky(2014) ; Carr(2020) . Many constraints can be placed on the mass of black holes that can serve as dark matter; the main open mass window lies in the asteroid mass range Carr(2021) ; Gorton(2024)
@@
Clearly, a sufficiently large PBH would have a significant impact on the human body; the goal of this present work is to determine the minimum black hole size needed to produce significant injury, and to see whether the observed lack of such injuries can place any bounds on PBH properties. Our discussion is partially motivated by earlier work on macroscopic dark matter (MACROs) Sidhu(2020) . In that paper, it was shown that the impact of MACRO dark matter on the human body would be sufficiently destructive that the non-observance of such impacts allows useful limits to be placed on the MACRO mass and cross section. While it is well known that a microscopic black hole would have a deleterious effect on the human body, there are no precise calculations. Most previous discussions of this topic are in the popular literature Loeb(2021) ; in fact, the first reference to the effect of a PBH on the human body appears to be in a work of fiction Niven(1974) . Lemos Lemos(1996) noted that very small black holes (less massive than those considered here) would have essentially no gravitational effect on the human body, but that the Hawking radiation could be quite destructive. Here we consider only gravitational forces, leaving the discussion of Hawking radiation for future work.